Follow this link to skip to                                      the main content

NASA News

Text Size

8 p.m. CDT, Thursday, Aug. 16, 2007
Mission Control Center, Houston, Texas
08.16.07
 
STATUS REPORT : STS-118-17
 
 
STS-118 MCC Status Report #17
 
 
Endeavour’s crew was informed just before bedtime today that mission managers had decided not to call upon them to repair the shuttle’s heat shield.

“Pass along our thanks for all the hard work the MMT (Mission Management Team) and everyone down there is doing supporting our flight,” said STS-118 Commander Scott Kelly.

As testing and deliberations related to the decision were under way, Kelly and Mission Specialists Rick Mastracchio, Dave Williams and Tracy Caldwell spent much of their day preparing for the possibility of a repair to a small ding in the silicon tiles on Endeavour’s belly.

Now, the fourth spacewalk of the mission on Saturday will remain focused on installation of two antennas and a stowage stand for the shuttle’s orbiter boom sensor system, and other tasks. Spacewalkers Williams and station Flight Engineer Clay Anderson also will retrieve two suitcase-sized packages known as Materials International Space Station Experiments 3 and 4. Those experiments were scheduled for retrieval at the end of Wednesday’s spacewalk, which was cut short due to a damaged spacesuit glove.

Also today, mission specialists Barbara R. Morgan and Alvin Drew answered questions from students at the Challenger Center for Space Science Education in Alexandria, Va. The duo then conducted interviews with Associated Press Television, Reuters and Idaho Public Television.

Morgan also answered questions via an amateur radio connection from students in her former school district, McCall-Donnelly in Idaho. Morgan taught at McCall-Donnelly Elementary School from 1975-1978, and again in 1979-1998.

Pilot Charlie Hobaugh and station Commander Fyodor Yurchikhin and Flight Engineer Oleg Kotov transferred supplies and equipment between the two orbiting spacecraft, which is more than half-way complete. Both crews closed the day with some off-duty time.

The next status report will be issued Friday morning or earlier if events warrant.